# The discrete periodic Pitman transform: invariances, braid relations, and Burke properties

### v2

@@ A.1. Basic algebraic properties
$=\left(\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(1-p_{2}q_{i})\right)p_{1}^{\mathfrak{s}(\widetilde{D}(\mathbf{k}_{2},\mathbf{k}_{1}))}p_{2}^{\mathfrak{s}(\widetilde{T}(\mathbf{k}_{2},\mathbf{k}_{1}))}\prod_{i=1}^{N}q_{i}^{\widetilde{D}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})+\widetilde{T}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}$
$=\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(p_{1}q_{i})^{\widetilde{D}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}\prod\limits_{i=1}^{N}(1-p_{2}q_{i})(p_{2}q_{i})^{\widetilde{T}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}$
$=\mathbb{P}(\mathbf{X}_{1}=\widetilde{D}(\mathbf{k}_{2},\mathbf{k}_{1}),\mathbf{X}_{2}=\widetilde{T}(\mathbf{k}_{2},\mathbf{k}_{1}))$
$=\mathbb{P}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{2},\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{1})$
where the first equality follows from the independence of $\mathbf{X}_{1},\mathbf{X}_{2}$ and $X_{m,i},X_{m,j}$ for $m\in\{1,2\}$ and $i,j\in\mathbb{Z}_{N}$ where $i\neq j$; the second equality follows from our definition of the geometric distribution; the third equality follows from Lemmas A.4(iii)-(iv) and A.5(i); and the final equality follows from the bijection in Lemma A.3. ∎
Proposition B.3 (Exponential Burke property).
@@
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})).$
@@
By Proposition B.2, we know that $(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}))$. Since $\max$, addition, and subtraction are continuous, $\widetilde{D}$ and $\widetilde{T}$ are continuous maps. We then have that
$\Bigl{(}\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\Bigr{)}\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}\Bigl{(}\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\Bigr{)}\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\mathbf{X}_{1},\mathbf{X}_{2}).\qed$
@@
The proof is the same as the proof of Theorems 1.7 and 1.8: replace the use of Theorem 1.4 with Theorem 1.10 and replace the Burke property in Propositions 4.1 with the Burke property in Propositions B.2 and B.3. ∎

### v3

@@ A.1. Additional algebraic properties
$=\left(\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(1-p_{2}q_{i})\right)p_{1}^{\mathfrak{s}(\widetilde{D}(\mathbf{k}_{1},\mathbf{k}_{2}))}p_{2}^{\mathfrak{s}(\widetilde{T}(\mathbf{k}_{1},\mathbf{k}_{2}))}\prod_{i=1}^{N}q_{i}^{\widetilde{D}_{i}(\mathbf{k}_{1},\mathbf{k}_{2})+\widetilde{T}_{i}(\mathbf{k}_{1},\mathbf{k}_{2})}$
$=\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(p_{1}q_{i})^{\widetilde{D}_{i}(\mathbf{k}_{1},\mathbf{k}_{2})}\prod\limits_{i=1}^{N}(1-p_{2}q_{i})(p_{2}q_{i})^{\widetilde{T}_{i}(\mathbf{k}_{1},\mathbf{k}_{2})}$
$=\mathbb{P}(\mathbf{X}_{1}=\widetilde{D}(\mathbf{k}_{1},\mathbf{k}_{2}),\mathbf{X}_{2}=\widetilde{T}(\mathbf{k}_{1},\mathbf{k}_{2}))$
$=\mathbb{P}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{1},\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{2})$
where the first equality follows from the independence of $\mathbf{X}_{1},\mathbf{X}_{2}$ and $X_{m,i},X_{m,j}$ for $m\in\{1,2\}$ and $i,j\in\mathbb{Z}_{N}$ where $i\neq j$; the second equality follows from our definition of the geometric distribution; the third equality follows from Lemmas 2.1 and 2.2(i); and the final equality follows from the bijection in Lemma A.5. ∎
Proposition A.7 (Exponential Burke property).
@@
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})).$
@@
By Proposition A.6, we know that $(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}))$. Since $\max$, addition, and subtraction are continuous, $\widetilde{D}$ and $\widetilde{T}$ are continuous maps. We then have that
$\Bigl(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})\Bigr)\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}\Bigl(\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\Bigr)\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\mathbf{X}_{1},\mathbf{X}_{2}).\qed$
@@
The proof is the same as the proof of Theorems 1.5 and 1.6: replace the use of Theorem 1.4 with Theorem 1.9 and replace the Burke property in Propositions 4.1 with the Burke property in Propositions A.6 and A.7. ∎